# Survey in Characterization of Semantic Change

### v2

@@ 1 Introduction
Language seems to be one of the most dynamic skills that humans have acquired through their evolution. It has constantly evolved to adapt the communication to new purposes and needs Pinker, (2003). While there is no single unified theory about its origins Allan, (2013), it is clear that language has undergone significant modifications since the early days of human history. These modifications have been driven by the need for humans to communicate complex ideas, interact with each other, pass on history and experiences, and build social bond McWhorter, (2011). One of the prevailing theories of language creation and evolution is based on the assumption of “continuity of change” Rilling et al., (2012). This theory holds that language is too complex to have arisen suddenly, and must have gradually evolved from earlier pre-linguistic systems used by our ancestors Pinker and Bloom, (1990). In this context, comprehending the process of language change is an contribution to the fields of history, sociology, linguistics, philosophy, and computer science ( Natural Language Processing (NLP)), as it is essential to understanding the origins of language, cultural development, grammatical theories, and the relation between language and thought Sherwood et al., (2008); Bowern, (2019).
@@
  • Phonological Changes changes in the creation, perception, and representation of sounds in language. For instance, the pronunciation of vowel sounds significantly changed due to the Great Vowel Shift in Medieval English Nevalainen and Traugott, (2016).
  • Lexical Changes alterations in the written form of words to form meaningful sentences and to pair with phonological changes, e.g., panem to pain, ‘bread’ in Latin to French.
  • Syntactic Changes alterations to the arrangement of words to create coherent sentences. For example, Modern English emphasizes word order and auxiliary verbs more than the intricate inflectional system used in Old English to identify grammatical connections Nevalainen and Traugott, (2016).
@@
While alterations in sound and form are crucial elements shaping our language, semantics plays a pivotal role in driving these changes Allan and Robinson, (2012). Semantic drift influences how words adapt to evolving cultural, technological, and societal contexts, creating new meanings, and preserving the language’s fundamental purpose of effective communication. Despite being a recurring topic in scientific research, semantic drifts remain less understood than other forms of linguistic evolution due to their often subtle and less observable nature, such as those influenced by grammar reforms.
@@
However, with the advent of computational linguistics and the development of advanced computer models, the field underwent a transformative shift. Linguists began harnessing the power of machine learning and natural language processing to automate the analysis of vast amounts of textual data. So the research community gradually shifted from anthropological observation to statistical methods for language analysis Allan and Robinson, (2012).
@@
These computer models, fueled by large corpora, and sophisticated algorithms, enabled researchers to detect subtle properties in language with unprecedented accuracy and efficiency. Enabling studying languages without any previous knowledge of it’s syntax or semantics, comparing multiple languages between themselves in the pursuit for universal properties Vulic et al., (2020).
@@
By embracing a model-based approach, linguistics can now unravel the intricacies of language phenomena and cope with the complexity of linguistics from different cultures and backgrounds Nivre et al., (2020). This scientific tooling provides a clear pathway for progress by facilitating direct comparisons between hypothesis, mathematical models, their refinement, and testing against emerging linguistic phenomena Hammarström et al., (2019).
While initially only observed with a theoretical motivation, the practical importance in computer science became very relevant. For example, in sentiment analysis, where the orientation of a word can change over time. Consider the word “sick”, while it traditionally meant “ill” or “unwell”, we can observe young people frequently using this word to mean “impressive” or “cool” Merriam-Webster, (2023). Another type of change is observed with the word “baby”. The frequent use of this word as a metaphor has impacted on its semantic interpretation in musics and poems, moving from “child” (as in the 14th century) to include the meaning “young woman” (sweetheart) in the 20th century.
The impact on applications like social media monitoring could be significant, as misinterpretation of slang terms, metaphors, or users’ domains could lead to unsuccessful outcomes. If an NLP algorithm is not aware about these differences, it may incorrectly interpret a sentence and attribute, for instance, a positive sentiment instead of a negative one (or vice versa), leading to an inaccurate analysis of the text Buechel et al., (2016).
@@
Addressing this issue involves continuously updating NLP models and algorithms to account for semantic change, incorporating real-time data sources, and staying aware of evolving language trends Lu et al., (2022). It underscores the need for adaptability, ongoing maintenance, and explain decisions of NLP applications that rely on natural language understanding.
The field of computational Lexical Semantic Change (LSC) emerged as more focused effort to improve methods to computationally identify semantic change in lexicons. Employing computer models that infer statistics from data has shown reliable capabilities to investigate lexicon meaning, enabling scholars to keep track of these variances in a theoretical and structured manner and many algorithms with diverse assumptions were proposed to improve meaning representation capabilities Tahmasebi et al., (2021), and consequently NLP systems.
@@
The development of better tools, that are cost effective, and can handle the many variation in language is crucial for many business. However, while the linguistic field has broadly studied what phenomena cause change Traugott, (2017), and many works provide deep analysis and theories on types of change. The computational linguistics community performed considerable effort to detect these drifts, nonetheless, to the best of our knowledge, less effort has been made to formalize and understand what kind of drift a word has undergone Hengchen et al., (2021).
@@
Another important development in this field is a proper language and formalism to describe the problems in semantic change Tahmasebi et al., (2021). Having a formal language is of paramount importance in a scientific field for several reasons. It provides a structured framework that allows researchers to precisely articulate and communicate their hypotheses and findings. It also ensures that ideas are expressed in a clear and unambiguous manner, facilitating collaboration and knowledge dissemination within the field. Furthermore, employing a mathematical language can significantly enhance our understanding of semantic change, advancing our comprehension of one of the most intriguing aspects of human communication.
The objective of this survey is to cover the lack of mapping and formalization in semantic change characterization. Being able to characterize such changes will open new perspectives in many areas. This is particularly the case in communication where semantic change creates potential for misunderstandings and ambiguity, especially in contexts where both meanings of a word are relevant, leading to a politically or socially misinterpretation of the context where the word was employed.
@@
As both computer science and linguistic communities, value understanding the computational aspects of language evolution. On this survey, we will focus on three factors of semantic changes: the orientation (i.e., if the words meaning is getting more pejorative or positive), the relation (i.e., if the words meaning change through a figurative usage), and the dimension (i.e., if the words meaning is becoming more abstract or more specific). Some of these properties, as we detail later, are crucial, for instance, for understanding and explaining the intention of a sentence, for writing tools to assist translation or to interact with users (e.g. in social robot). Finally, it will open new opportunities to improve reasoning capabilities for NLP tools Danilevsky et al., (2020).
Drawing from a comprehensive review of the extensive literature across related areas, this survey makes a significant contribution by consolidating, categorizing, and pinpointing gaps in existing approaches. Furthermore, it introduces formal descriptions and proposes enhancements towards a more robust characterization of semantic change. This endeavor represents a pivotal advancement in our comprehension of this intricate phenomenon. By synthesizing the collective knowledge from prior research and offering systematic solutions to the identified limitations, this survey not only helps streamline the existing body of work but also paves the way for a more efficient and comprehensive understanding of the multifaceted nature of semantic change.
@@
The remainder of the paper is structured as follows: Section 2 reviews existing surveys and their difference with respect to this work. In Section 3 we investigate available corpora of documents for semantic change identification, characterization and their limitations. In Section 4 we explore existing methods to represent meaning change characterization and how they attempt to solve the characterization problem. Section 6 presents our analysis and open questions in characterization. Finally, Section 7 we conclude this survey and outline future work.
@@ 2 Lexical Semantic Change
Concerning the characterization of change, we adopt in this survey the mainstream typology in linguistic literature for semantic change Koch, (2016); Blank et al., (2003); Hock and Joseph, (2019), (i) broadening and narrowing, (ii) amelioration and pejoration, and (iii) metonymization and metaphorization. Before starting our study on change characterization, we reviewed the works published on change identification. This is a well-developed topic that focuses on evaluating whether the meaning of a word has changed or not. Papers addressing change identification often claim the need for further understand and characterization of the type of change that occurred, but the current state of works on computational change characterization is unclear. This conclusion motivated the work presented in this paper and differentiate our survey from existing ones on LSC.
@@ 2.1 Related Works on Identification of Lexical Semantic Changes
Surveys for semantic change identification are regularly published, with a main focus on methodologies for change detection. In a survey about “semantic change computation” Tang, (2018), Tang proposes to split LSC analysis into 4 steps: obtaining and preparing a diachronic corpus, identifying word senses, modeling semantic change, followed by presentation and validation of hypotheses. This framework sets the structure of the survey, describing the studies related to each of these steps. The characterization of changes is mainly discussed in the “modeling semantic change” section, where the author explains the difference between Cambridge Change Ardley, (1968), and McTaggart Change Broad, (1938) and cites the approaches following these definitions. However, the approaches that can classify semantic changes as sense broadening, sense narrowing, pejoration, amelioration, metaphorical change and metonymic change are not included.
The survey about computational approaches to LSC Tahmasebi et al., (2018) focuses on word-level differentiation, sense differentiation, and lexical replacement. While it includes approaches with several types of change categories (e.g., metaphor, metonymy), the main focus of change categories are broadening, narrowing and correlated types (split, join, birth). However, one recommendation of the authors is to have deepen studies on methods to automatically detecting other types of change and to use it to build the history of the words. Auxiliary to the methodologies discussed, the authors provide an in-depth analysis of linguistic literature that grounds each domain, followed by evaluation strategies and applications. The difficulty to compare approaches is highlighted in the survey and, according to the authors, is the result of an absence of standard metrics/practices and reference datasets.
@@
First studies on the field employed graph- or frequency-based methods to analyse LSC, later approaches benefited from the advances in the state-of-the-art to replace the analytical methods by those based on embeddings and contextual embeddings. Although contextual embeddings represented a huge improvement for a series of NLP tasks Devlin et al., (2019), the improvement for LSC did not follow this trend Schlechtweg et al., (2020), showing that there is still room for improvement in this domain Zhou et al., (2021); Kutuzov et al., (2022).
@@
Research in this field also delves into ongoing LSC methodological issues. These issues involve the challenges of characterizing semantic shifts. As pointed out in Hengchen et al., (2021), there are few studies that detail the nature of these changes, mainly because it is difficult to model meaning from textual data. It further concludes that no particular measures for quantifying category change exist and that a consensus on taxonomy for classifying types of change is required.
@@
For this survey, we argue that semantic change characterization needs to have appropriated methods and tools to analyse each fine-grained category. We also defend that observing variations in a single property can not describe well all forms of LSC. For instance, changes in angles between vectors representing words Dubossarsky et al., (2017) is not the best solution to identify narrowing/broadening of sense Tahmasebi et al., (2018). Specific methods for identifying the different types of change could led to a better understanding of LSC than straightforward dissimilarity measures.
@@ 2.2 Objectives
Different terms are used to refer to Lexical Semantic Change (LSC), such as semantic drift or semantic shift. In this paper we make the distinction between identification of a change and characterization of the change. The former refers to the process through which the meaning of a word gradually or significantly changes over time in a language. It is the process through which a language unit’s initial semantic content changes, leading to a shift in its usage, connotation, or meaning. Cambridge Ardley, (1968) and McTaggart Broad, (1938) change approaches are different on how to use temporal information to identify changes. While Cambridge proposes to compare two different corpus (from different time period), McTaggart prefers to study the evolution of the usage of a word over a period of time. The analysis of works dealing with the identification problem has been well covered by Tahmasebi et al., (2018), Kutuzov et al., (2018), and Tang, (2018). We formally define the semantic change identification in Section 5.1.
The focus of our literature review is on the characterization of change, where we look for common properties that allows for grouping identified semantic changes into categories. Semantic change characterization occurs by the need to reduce fuzziness when combining meanings and forms Hock and Joseph, (2019). While some authors propose to use the cause of the changes to group them, other authors categorize the type of change by language figures that describe the change, or by the cultural process. From a computation perspective, determining the cause of a semantic change becomes very complex to implement and, to the best of our knowledge, none computational approach was implemented yet. For this review we followed the proposition of using language figures as bases to create categories of changes. We adopt the most predominant categories in literature, following the typology proposed in Juvonen and Koptjevskaja-Tamm, (2016) and Tahmasebi et al., (2021):
@@
  • Narrowing restriction of meaning where a symbol represents fewer concepts, as in ‘gay’ that used to have senses as festive or happy, now it generally is used in a single sense of homossexual.
@@
  • Metaphorization conceptualizing one thing in terms of another, e.g., the use of Latin ad ’to’ + mit ’send’ for locution (admit) Traugott, (2017).
As the types in the above typology are directly correlated (involves the same computational challenge), e.g., broadening-narrowing, requires sense counting, amelioration-pejoration, requires sentiment analysis, metaphorization-metonymization, requires some form of relatedness measure Koch, (2016); Blank et al., (2003); Hock and Joseph, (2019), we group them in poles to better compare, categorize, and generalize studies exploring these types of change. Nevertheless, we have included only approaches that consider changes between corpora as this typology was designed for diachronic studies, not changes related to a specific domain or to a specific context Koch, (2016); Blank et al., (2003); Hock and Joseph, (2019). In Figure 1 we illustrate the hierachical taxonomy for LSC.
@@
We assign studies that track sense broadening or narrowing in the dimension ($\mathcal{D}$) pole and we assume that words that are getting more or fewer senses are increasing or decreasing their dimension value (see Section 5.2). This pole is self-complementary as the change in the value of the dimension addresses both senses. A change in dimension could work as a level of carefulness with translations for example, that broader words should be translated with more care Mohammed, (2009).
We regroup in the relation ($\mathcal{R}$) pole changes increasing the relation between one word sense and other senses for this word (see Section 5.3). This kind of changes are commonly related to metaphorization or metonymization, where the meaning of a word depends on the link built by a conceptual (abstract) or material (direct) similarity between concepts. Relations like hypernymy, homonymy, and synonymy are also factors for semantic change Koch, (2016); Hock and Joseph, (2019) as they are material relation and can potentially be encompassed in this pole.
For the orientation ($\mathcal{O}$) pole, we group studies that identify the process of amelioration or pejoration or any perceptual change in the ”sentiment” related to words (e.g., criticism, compliment, etc.). This pole also encompasses many forms of change in perception of opinions, as some studies suggest that sentiments like ‘taboo’ Koch, (2016) also affect the change in word usage.

### v3

@@ 1 Introduction
Language seems to be one of the most dynamic skills that humans have acquired through their evolution. It has constantly evolved to adapt the communication to new purposes and needs (Pinker,, 2003, p. 13-23). While there is no single unified theory about its origins (Allan,, 2013, p. 13-52), it is clear that language has undergone significant modifications since the early days of human history. One of the prevailing theories of language creation and evolution is based on the assumption of “continuity of change” (Rilling et al.,, 2012). This theory holds that language is too complex to have arisen suddenly, and must have gradually evolved from earlier pre-linguistic systems used by our ancestors (Pinker and Bloom,, 1990, p. 725-726). In this context, comprehending the process of language change is an contribution to the fields of history, sociology, linguistics, philosophy, and Natural Language Processing (NLP) (Campbell,, 2013, p. 1-8), as it is essential to understanding the origins of language, cultural development, grammatical theories, and the relation between language and thought (Sherwood et al.,, 2008, p. 426–454).
@@
  • Phonological Changes changes in the creation, perception, and representation of sounds in language. For instance, the pronunciation of vowel sounds significantly changed due to the Great Vowel Shift in Medieval English (Nevalainen and Traugott,, 2016, p. 367).
  • Orthographic Changes alterations in the written form of words to form meaningful sentences and to pair with phonological changes, e.g., panem to pain, ‘bread’ in Latin to French.
  • Syntactic Changes alterations to the arrangement of words to create coherent sentences. For example, Modern English emphasizes word order and auxiliary verbs more than the intricate inflectional system used in Old English to identify grammatical connections (Nevalainen and Traugott,, 2016, p. 616).
@@
While alterations in sound and form are crucial elements shaping our language, semantics plays a pivotal role in driving these changes. Semantic drift influences how words adapt to evolving cultural, technological, and societal contexts, creating new meanings, and preserving the language’s fundamental purpose of effective communication. Despite being a recurring topic in scientific research, semantic drifts remain less understood than other forms of linguistic evolution due to their often subtle and less observable nature, such as those influenced by grammar reforms (Allan and Robinson,, 2012, p. 2-3).
@@
However, with the advent of computational linguistics and the development of advanced computer models, the field underwent a transformative shift. Linguists began harnessing the power of machine learning and natural language processing to automate the analysis of vast amounts of textual data. So the research community gradually shifted from anthropological observation to statistical methods for language analysis (Allan and Robinson,, 2012, p. 161-164).
@@
These computer models, fueled by large corpora, and sophisticated algorithms, enabled researchers to detect subtle properties in language with unprecedented accuracy and efficiency. Enabling studying languages without any previous knowledge of it’s syntax or semantics, comparing multiple languages between themselves in the pursuit for universal properties (Vulic et al.,, 2020).
@@
By embracing a model-based approach, linguistics can now unravel the intricacies of language phenomena and cope with the complexity of linguistics from different cultures and backgrounds (Nivre et al.,, 2020). This scientific tooling provides a clear pathway for progress by facilitating direct comparisons between hypothesis, mathematical models, their refinement, and testing against emerging linguistic phenomena (Hammarström et al.,, 2019).
While initially only observed with a theoretical motivation, the practical importance in computer science became very relevant. Historical knowledge graphs can contain thousands of entities and relations that change over time (e.g., Wikidata and English WordNet) – automatically identifying concepts that change can allow an automatic update of these systems and enable semantic resources to extend the taxonomy of preexisting concepts (Polleres et al.,, 2023).
Other examples of applications that are impacted by semantic change:
  • Personalized vernacular voice assistants. In Brazil, ‘bicha’ can mean “homosexual,” “unidentified female,” or “queue,” depending on which state this assistant is located in. Therefore, it is desirable for an agent using vernacular communication to be aware of this distinction.
  • Region-based translation. In French, ‘déjeuner’ can mean “breakfast,” “lunch,” or “dinner,” depending on which French-speaking country you are addressing the translation. So, a correct translation should respect the geographic origin of the text.
  • Social network analysis. Some political labels have different meanings and orientations depending on which sub-network you analyze (e.g., “We need to change. MAGA!” meaning the political slogan, with a positive sense or “He is a MAGA” meaning he support a politician who uses this slogan, with a negative sense). So, under-specified sentences could mislead sentiment analysis systems.
  • Merging domain ontologies. The word ‘machine’ can mean “a computer program” in computer science literature or “a mechanical device” in engineering literature. So, the entity ‘machine’ does not hold all the relations between both ontologies.
@@
Addressing this issue involves continuously updating NLP models and algorithms to account for semantic change, incorporating real-time data sources, and staying aware of evolving language trends (Lu et al.,, 2022). It underscores the need for adaptability, ongoing maintenance and engineering, and explain decisions of NLP applications that rely on natural language understanding. Specially in the case of Language Models where regular updates requires substantial computational resources, time, and financial investment.
@@
The development of cost-effective tools that can handle the many variations in language and precisely understand the change in semantics is crucial for many businesses. However, while the linguistic field has broadly studied what phenomena cause change (Traugott,, 2017), and many works provide deep analysis and theories on types of change. The computational linguistics community performed considerable effort to detect these drifts, nonetheless, to the best of our knowledge, less effort has been made to formalize and understand what kind of drift a word has undergone (Hengchen et al.,, 2021).
@@
Another important development in this field is a proper language and formalism to describe the problems in semantic change (Kutuzov et al.,, 2018). Having a formal language (i.e., a mathematical description) is of paramount importance in a scientific field for several reasons. It provides a structured framework that allows researchers to precisely articulate and communicate their hypotheses and findings. It also ensures that ideas are expressed in a clear and unambiguous manner, facilitating collaboration and knowledge dissemination within the field. Furthermore, employing a mathematical language can significantly enhance computational interpretation of semantic change, advancing our comprehension of one of the most intriguing aspects of human communication.
This survey addresses the need to understand and formalize semantic change’s characterization. Characterizing such changes will open new perspectives in various areas, mainly because we have no resources to study them [^1] While NLP resources have hundreds of thousands of examples, the datasets of semantic change generally cover less than a hundred words Kutuzov et al., (2022).. This is especially problematic in contexts where both word meanings are relevant, potentially leading to political or social misinterpretations. The survey’s objectives include:
  1. 1.
    Identifying the limitations of current representations and methodologies in the categorization of semantic changes.
  2. 2.
    Analyzing the definitions of each category and the complexity of computer algorithms’ interpretations.
  3. 3.
    Evaluating the strengths of approaches used to classify lexical semantic changes (LSC).
  4. 4.
    Finding and analyzing formal definitions based on first-order logic to reduce ambiguity and facilitate the comparison of findings.
@@
There are many distinct categories for semantic change, ranging from cultural tradition (e.g., taboo) to cognitive aspects (e.g., metaphor). On this survey we select categories that are both addressable by computational methods on a comparative setting (Tang et al.,, 2013) and by linguistic literature (Campbell,, 2013, p. 221-246), i.e., changes we can systematically measure and are backed by a theoretical analysis (Hammarström et al.,, 2019). We will focus on three factors of semantic changes: the orientation (i.e., if the words meaning is getting more pejorative or positive), the relation (i.e., if the words meaning change through a figurative usage), and the dimension (i.e., if the words meaning is becoming more abstract or more specific). Some of these properties, as we detail later, are crucial, for instance, for understanding and explaining the intention of a sentence, for writing tools to assist translation or to interact with users (e.g. in social robot). Finally, it will open new opportunities to improve reasoning capabilities for NLP tools (Danilevsky et al.,, 2020).
This survey is the first to comprehensively cover semantic change characterization by consolidating and categorizing existing approaches and identifying gaps. It introduces formal descriptions and proposes enhancements for a more robust understanding of semantic change. By synthesizing prior research and addressing its limitations, the survey streamlines existing work and advances our understanding of this complex phenomenon.
@@
The remainder of the paper is structured as follows: Section 2 reviews existing surveys and their difference with respect to this work. In Section 3 we investigate available corpora of documents for semantic change identification, characterization and their limitations. In Section 4 we explore existing methods to represent meaning change characterization and how they attempt to solve the characterization problem. Section 5 we present a formal model derived from reviewed methodologies and linguistic description. Section 6 presents our analysis and open questions in characterization. Finally, Section 7 we conclude this survey and outline future work.
@@ 2 Lexical Semantic Change
Concerning the characterization of change, we adopt in this survey the mainstream typology in linguistic literature for semantic change (Traugott,, 2017), (i) broadening and narrowing, (ii) amelioration and pejoration, and (iii) metonymization and metaphorization. Before starting our study on change characterization, we reviewed the works published on change identification. This is a well-developed topic that focuses on evaluating whether the meaning of a word has changed or not. Papers addressing change identification often claim the need for further understand and characterization of the type of change that occurred, but the current state of works on computational change characterization is unclear. This conclusion motivated the work presented in this paper and differentiate our survey from existing ones on Lexical Semantic Change (LSC).
@@ 2.1 Related Works on Identification of Lexical Semantic Changes
Surveys for semantic change identification are regularly published, with a main focus on methodologies for change detection. In a survey about “semantic change computation” (Tang,, 2018), Tang proposes to split LSC analysis into 4 steps: obtaining and preparing a diachronic corpus, identifying word senses, modeling semantic change, followed by presentation and validation of hypotheses. This framework sets the structure of the survey, describing the studies related to each of these steps. The characterization of changes is mainly discussed in the “modeling semantic change” section, where the author explains the difference between Cambridge Change (Ardley,, 1968, p. 91), and McTaggart Change (Broad,, 1938, p. 297) and cites the approaches following these definitions. However, the approaches that can classify semantic changes as sense broadening, sense narrowing, pejoration, amelioration, metaphorical change and metonymical change are not included.
The survey about computational approaches to LSC (Tahmasebi et al.,, 2018) focuses on word-level differentiation, sense differentiation, and lexical replacement. While it includes approaches with several types of change categories (e.g., metaphor, metonymy), the main focus of change categories are broadening, narrowing and correlated types (split, join, birth). However, one recommendation of the authors is to have deepen studies on methods to automatically detecting other types of change and to use it to build the history of the words. Auxiliary to the methodologies discussed, the authors provide an in-depth analysis of linguistic literature that grounds each domain, followed by evaluation strategies and applications. The difficulty to compare approaches is highlighted in the survey and, according to the authors, is the result of an absence of standard metrics/practices and reference datasets.
@@
First studies on the field employed graph- or frequency-based methods to analyse LSC, later approaches benefited from the advances in the state-of-the-art to replace the analytical methods by those based on embeddings and contextual embeddings. Although contextual embeddings represented a huge improvement for a series of NLP tasks (Devlin et al.,, 2019), the improvement for LSC did not follow this trend (Schlechtweg et al.,, 2020), showing that there is still room for improvement in this domain (Zhou et al.,, 2021; Kutuzov et al.,, 2022).
@@
Research in this field also delves into ongoing LSC methodological issues. These issues involve the challenges of characterizing semantic shifts. As pointed out by Hengchen et al., (2021), there are few studies that detail the nature of these changes, mainly because it is difficult to model meaning from textual data. It further concludes that no particular measures for quantifying category change exist and that a consensus on taxonomy for classifying types of change is required.
@@
For this survey, we argue that semantic change characterization needs to have appropriated methods and tools to analyse each fine-grained category. We also defend that observing variations in a single property can not describe well all forms of LSC. For instance, changes in angles between vectors representing words (Dubossarsky et al.,, 2017) is not the best solution to identify narrowing/broadening of sense (Tahmasebi et al.,, 2018). Specific methods for identifying the different types of change could lead to a better understanding of LSC than straightforward dissimilarity measures.
@@ 2.2 Objectives
Different terms are used to refer to Lexical Semantic Change (LSC), such as semantic drift or semantic shift. In this paper we make the distinction between identification of a change and characterization of the change. The former refers to the process through which the meaning of a word gradually or significantly changes over time in a language. While the following refers to which characteristic could describe and differentiate the process that lead to a change in meaning. We formally differentiate this setting in Section 5.1. The analysis of works dealing with the identification problem has been well covered by Tahmasebi et al., (2018), Kutuzov et al., (2018), and Tang, (2018).
Semantic change occurs by the need to reduce fuzziness when combining meanings and forms (Hock and Joseph,, 2019). While some authors propose to use the cause of the changes to group them, other authors categorize the type of change by language figures that describe the change, or by the cultural process. The focus of our literature review is on the characterization of change, where we look for common properties that allows for grouping identified semantic changes into categories. From a computation perspective, determining the cause of a semantic change becomes very complex to implement and, to the best of our knowledge, none computational approach was implemented yet.
Some frameworks were proposed on how to pose and track LSC like Cambridge and McTaggart (Tang et al.,, 2013), which initially were not considering computational methods to it. Cambridge (Ardley,, 1968) and McTaggart (Broad,, 1938) change approaches are different on how to use temporal information to identify changes, while Cambridge proposes to compare two different corpus (from different time period), McTaggart prefers to study the evolution of the usage of a word over a period of time. The approach that best suits the computational perspective is the Cambridge, as McTaggart approach would require historiographer work to acquire historic socio-cultural knowledge that goes beyond contextual analysis of word occurrence (Tang et al.,, 2013, p. 4-7).
For this review we assume the Cambridge setting for change analysis using language figures as bases to create categories of changes, which align with current state-of-the-art methods for computational LSC (Pivovarova and Kutuzov,, 2021). Given this considerations, we adopt the Traugott, (2017) typology, which is also the most predominant categories in literature, as in Koch, (2016, p. 21-66), Hock and Joseph, (2019, p. 189-222), and Campbell, (2013, p. 221-246):
@@
  • Narrowing restriction of meaning where a symbol represents fewer concepts, as in ‘gay’ that used to have senses as festive or happy, now it generally is used in a single sense of homosexual.
@@
  • Metaphorization conceptualizing one thing in terms of another, e.g., the use of Latin ad ’to’ + mit ’send’ for locution (admit) (Traugott,, 2017).
As the types in the above typology are directly correlated (involves the same computational challenge), e.g., broadening-narrowing, requires sense counting, amelioration-pejoration, requires sentiment analysis, metaphorization-metonymization, requires some form of relatedness measure (Koch,, 2016; Blank et al.,, 2003; Hock and Joseph,, 2019), we group them in poles to better compare, categorize, and generalize studies exploring these types of change. Nevertheless, we have included only approaches that consider changes between corpora as this typology was designed for diachronic studies, not changes related to a specific domain or to a specific context (Koch,, 2016; Blank et al.,, 2003; Hock and Joseph,, 2019). In Figure 1 we illustrate the hierachical taxonomy for LSC.
@@
We assign studies that track sense broadening or narrowing in the dimension ($\mathcal{D}$) pole and we assume that words that are getting more or fewer senses are increasing or decreasing their dimension value (see Section 5.2). This pole is self-complementary as the change in the value of the dimension addresses both senses. A change in dimension could work as a level of carefulness with translations for example, that broader words should be translated with more care (Mohammed,, 2009).
We regroup in the relation ($\mathcal{R}$) pole changes increasing the relation between one word sense and other senses for this word (see Section 5.3). This kind of changes are commonly related to metaphorization or metonymization, where the meaning of a word depends on the link built by a conceptual (abstract) or material (direct) similarity between concepts. Relations like hypernymy, homonymy, and synonymy are also factors for semantic change (Koch,, 2016; Hock and Joseph,, 2019) as they involve a material relation (related by some cultural association) and can potentially be encompassed in this pole.
For the orientation ($\mathcal{O}$) pole, we group studies that identify the process of amelioration or pejoration or any perceptual change in the ”sentiment” related to words (e.g., criticism, compliment, etc.). This pole also encompasses many forms of change in perception of opinions, as some studies suggest that sentiments like ‘taboo’ (Koch,, 2016) also affect the change in word usage.